OpenStudy (anonymous):

Find the exact value, in radians, of the expression. sin -1(√2/2)

jimthompson5910 (jim_thompson5910):

Hint: use a unit circle http://www.math.ucsd.edu/~jarmel/math4c/Unit_Circle_Angles.png

OpenStudy (anonymous):

How do i tell which side is the sin?

jimthompson5910 (jim_thompson5910):

the y coordinate of any point on the unit circle is equal to sin(theta)

jimthompson5910 (jim_thompson5910):

example: sin(60) = sqrt(3)/2 because the point at 60 degrees has sqrt(3)/2 as the y coordinate sin(pi/3) = sqrt(3)/2 ... same reason as above (just in the angle is now in radians)

OpenStudy (anonymous):

\[5\pi/4\]?

OpenStudy (anonymous):

Ok...thanks thats makes sense!!

jimthompson5910 (jim_thompson5910):

there's one before that in quadrant 1

jimthompson5910 (jim_thompson5910):

what you see is in quadrant 2

jimthompson5910 (jim_thompson5910):

technically you are correct, but they want the principal value

jimthompson5910 (jim_thompson5910):

oh wait, you said 5pi/4 when it should be 3pi/4

OpenStudy (anonymous):

yes! Thanks

jimthompson5910 (jim_thompson5910):

anyways, take a look at the angle and point I circled You'll see the answer is actually pi/4

OpenStudy (anonymous):

So if tan^-1 0 would be 3pi/2?

jimthompson5910 (jim_thompson5910):

no tan is undefined at pi/2 and 3pi/2

jimthompson5910 (jim_thompson5910):

but tan is zero when theta = 0 or theta = pi so that's why tan^-1(0) = 0

OpenStudy (anonymous):

Interesting...you explain this so much easier than my professor

jimthompson5910 (jim_thompson5910):

glad you're understanding it